Sally travels 990m south 486m west find the displacement and distance?
1
Expert's answer
2020-03-12T02:50:19-0400

The distance


"d=990+486=1476\:\rm m."

The displacement


"s=\sqrt{990^2+486^2}=1103\:\rm m."